Jazz World Inc. is considering a project that has the following cash flow and WACC data. What is the project's NPV?
 

WACC:

12.00%





Year

0    

1  

2  

3  

4  

Cash flows

-$1,200

$400

$425

$450

$475
Answer #1

K47 fc =SUM(K42:46) G J K 40 H Calculation of NPV Year 41 42 0 43 1 Cashflows (1,200.00) 400.00 425.00 450.00 475.00 44 $ $ $K47 G K 40 41 0 42 43 fc =SUM(K42:K46) H J Calculation of NPV Year Cashflows Pv@ 12% -1200 =1 1 400 =J42/1.12 425 =J43/1.12 3